what is the slope of the line y=3x
Answer by stanbon(48535)

what is the slope of the line y=3x
-----------------
Generate a couple of points and calculate the slope.
If x = 0, y =3*0=0 giving the point (0,0)
--------
If x = 1, y=3*1=3 giving the point (1,3)
-----------
Calculate the slope:
slope = [3-0]/[1-0] = 3/1 = 3
============
Cheers,
Stan H.